solve for real roots: sqrt(2x^2+x-12)=x
this is not a text book, just on paper, I do not understand it

Answer by stanbon(75887)

sqrt(2x^2+x-12)=x
Square both sides to get:
2x^2+x-12 = x^2
x^2+x-12 = 0
(x+4)(x-3)=0
x= -4 or x=3
Checking in the original equation:
If x=-4: sqrt(32-4-12)=sqrt16 = -4
No; that is wrong so x=-4 is not a solution.
----------
sqrt(2*9+3-12)= sqrt(9)=3
Right; x=3 is a solution.
===========
cheers,
Stan H.